Suppose a firm uses two variable inputs: L and K. Initially employs K=15 units and L=5units; and produced, Q=3005 units denoted by an Isoquant curve. Assuming Marginal Product of L, MPL=6 units and MPK=18, answer the following question.
A. Determine the economically efficiant level of labor and captital that maximizes profit.

B. If the firm reduces the eployment by 10, by how much L has to be increased to keep the level of output constant on the isoquant.
1
Expert's answer
2020-01-08T09:33:09-0500

A. At the economically efficiant level of labor and captital MPL/w = MPK/r and w*L + r*K = Q, so:

6/w = 18/r,

r = 3w,

w*5 + 3w*15 = 3005,

50w = 3005,

w = 60.1, r = 180.3.

B. If the firm reduces the employment by 10, then to keep the level of output constant on the isoquant, then:

60.1*L + 180.3*(15 - 10) = 3005,

L = 35.